Simplify (cuberoot (sqrt ( 245 ) - sqrt ( 20 ) ) )^(-2)
Answer by stanbon(75887)

Simplify (cuberoot (sqrt ( 245 ) - sqrt ( 20 ) ) )^(-2)
-----
(cbrt(7sqrt(5)-2sqrt(5))^(-2)
-----
= cbrt(5sqrt(5))^-2
= cbrt (5^(3/2))^-2
= (5^(1/2)^(-2)
= 5^-1
= 1/5
-------------
Cheers,
Stan H.
--------------